OpenStudy (anonymous):

if x=-3 what is the value of 2x^2-2x+1

OpenStudy (anonymous):

To solve this, we substitute the given value of x (x = -3) in 2x^2-2x+1.

OpenStudy (anonymous):

Therefore: 2(-3)^2 - 2(-3) + 1 = 2(9) + 6 +1 = 18 + 6 + 1 = ?

OpenStudy (anonymous):

choices are -11,1,7,190r -19

OpenStudy (anonymous):

Ah. Then I'm lost. Sorry :/

OpenStudy (anonymous):

I will answer this!

OpenStudy (anonymous):

It will come out to be -31

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

is everything behind the ^ a power?

OpenStudy (anonymous):

nevermind, I can't answer this...

OpenStudy (solomonzelman):

@embeth3719 2x^2-2x+1 plug in -3 2(-3)^2-2(-3)+1 18+6+1 25 Your choices are wrong, this is the correct answer.

OpenStudy (solomonzelman):

@Titanium.022, you were correct!

OpenStudy (anonymous):

I thought I was... but I wasn't sure. xD

OpenStudy (solomonzelman):

Yeah, I see. For sure correct! Sometimes the choices are wrong! @embeth3719, are you sure this is the Correct question, and this a the correct choices? Perhaps you missed something while typing it?